PROBLEM:
The Vang family has a backyard with the dimensions of 35ft, 75ft, 90ft. They want to plant small trees that require 9 square feet of surface area each. How many should they plant? Round your answer to the nearest whole tree.





Answer by checkley77(12569)

Because this is not a right triangle.
35^2+75^2 does not = 90^2
1225+5625 does not = 8100
6850 does not = 8100
The folowing formula for area is used:
A=SQRT(S-A)(S-B)(S-C)
WHERE S=(A+B+C)/2=(35+75+90)/2=200/2=100
A=SQRT(100-35)(100-75)(100-90)
A=SQRT(65*25*10)
A=SQRT16250
A=127 FT^2
127/9=14 TREES CAN BE PLANTED.